Solve each literal equation for the indicated variable
A=1/2 h(B+b) (for b) area of trapezoid fraction

Answer by checkley75(3666)

A=1/2 H(B+b)
A=H(B+b)/2
2A=HB+Hb
Hb=2A-HB
b=(2A-HB)/H